OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x - 2y = 2

OpenStudy (anonymous):

Substitution: Find the value of y in the first equation: 4x + y = 8 y = -4x + 8 Now SUBSTITUTE the value of y into the second equation: x - 2y = 2 x - 2(-4x + 8) = 2 x + 8x - 16 = 2 9x - 16 = 2 Can you solve for x from there?

OpenStudy (anonymous):

YEP! lol